Previous

Calendar

June

2007

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

27

JD147

28

JD148
CU Holiday (Memorial Day)

29

JD149
between 11-12:30 MDT, an intense storm passed by the tower

the flow rate (ie, pump) for the LI-6262 was acting strange (on and off thoughout the day).

30

JD150
Site Visit (SpB, JB): Details , Photos (27)

at 12:30 MST, replaced regulator for span gas at tower

from ~11:30-12:30 MST, replaced ec gast pump with refurbished pump; made flow adjustments at LI-6262

at 11:50 MST, started running "met_02.csi" on MET logger (output RM Young in mV to quacker)

31

JD151 (blustery)
Site Visit (SpB, JB): Details , Photos (15)

from ~11:00-15:30 MST, replaced gelman inlet filters on profile system (1,2,6,10, and 21.5m)

from 10:52-12:30 MST, running ops41 (removed handar.9m & modified 2-d ati code)

at 12:37 MST, start running ops42 (removed 2-d ati, added back NCTC's)

Swapped 4 flasks at west robot (appears to be pump problems?)

at 10:45 MST, started running "met_03.csi" on MET logger (removed 9m handar output)

1

JD152

2

JD153

3

JD154

4

JD155

5

JD156

6

JD157

7

JD158
Site Visit (JH, ?): Details

very windy (night of Jun 6th and morning/day of Jun 7th)

8

JD159
Site Visit (SpB, JB, NT, SL): Details , Photos (27)

from ~10:30-11:00 MST, swapped out gast pump for LI-6262 (replaced with a new pump)

2-d ati sonic was having "cross-talk" problems

Pump at west robot seems ok, switched to 15-flask "summer" mode (AUTO26NW.CSI); no flasks swapped

9

JD160

10

JD161

11

JD162 (cool; light rain)
Site Visit (SpB, JB, JH, SL, MP): Details , Photos (6)

LTER site review

12

JD163

13

JD164

14

JD165 (warm, clear)
Site Visit (JB, SpB): Details , Photos (19)

Flag Day

Added new trh sensor to channel 204 (deployed at met logger); started running ops43

Modified met logger program for new trh sensor; started running MET_04.DLD

JB swapped 4 flasks and moved lowest inlet to ~10cm above ground

SpB replaced battery for 23xsoil logger

MRS logger data downloaded, and logger dis-connected from power (for summer).

russter2 rebooted (OS update)

15

JD166
SpB Off (Start)
sometime between 15:00-16:00 MDT, Wyoming King Air passes over the CU tower (DM)

16

JD167

17

JD168
Fathers Day

18

JD169
Site Visit (JH,MP,SLS,DK) :
Details


19

JD170 (warm, clear)
Site Visit (JB,MP,SL,SLS,LL): Details , Photos (7)

Swapped edgeports at quacker; still running ops43

JB took forest phenol photos

JB removed inlet filters and boxes from hydra's tubing

20

JD171
Site Visit (JB): Details

Total power loss to tower and trailer; data system down ~3:45PM

JB went up to site after power loss to check things out and shutdown quacker to avoid battery burn; could not determine cause of power loss

21

JD172 (sunny, warm)
Site Visit (JB,SL,RM,RKM,LL): Details

Summer Solstice

Power out all day

JB, Jielun, and RKM organized and layed out HYDRA's dekabon tubing @ Cuomo Creek

22

JD173

Site Visit (JB,JH,SLS): Details

Power still out as of A.M.

power restored ~2 P.M.; JB rebooted quacker, swapped span cylinders, swapped 15 flasks at robot

23

JD174
Site Visit (LL): Details

24

JD175
Site Visit (LL): Details

25

JD176 (warm, dry)
Site Visit (JB, SpB, LL): Details , Photos (16)

Downloaded from soil and Tcs loggers; moved prof. pump to second galvanized "power box" on cuff tower

SpB took forest phenol photos

SpB Returns

26

JD177 (foggy)
Site Visit (SpB, JH, SL, MP, SS): Details

started to work on HYDRA (using thinkpad)

setout new litter chambers (MP/SL/JH)

27

JD178

28

JD179

29

JD180
Site Visit (JB,SpB,LL,RaM):
Details
, Photos (2)

JB cleaned KrH2O optics at 11:20 MST

JB swapped 15 flasks

30

JD181